please help me out...........

-54=10-(m-10)^3/2

Answer by stanbon(57299)

-54=10-(m-10)^3/2
---
(m - 10)^(3/2) = 64
---
Raise both sides to the 2/3 power to get:
m - 10 = 64^(2/3)
--------
Take the cube root of the right side:
m-10 = 4^2
Square the right side
m - 10 = 16
m = 26
============
Cheers,
Stan H.